OpenStudy (anonymous):

If 7p + 5q = 2 and 8p - 9q = 17, then -p + 14q =

jimthompson5910 (jim_thompson5910):

Hint: (Equation 1) - (Equation 2)

OpenStudy (anonymous):

either multiply the first by 8 and the second by 7 and subtract the two new equations or multiply the first by 9 and the second by 5 and add the two new equations

jimthompson5910 (jim_thompson5910):

there's no need to solve for p and q individually, but you can if you want

OpenStudy (texaschic101):

7p + 5q = 2 -->(9)7p + 5q = 2 8p - 9q = 17 -->(5)8p - 9q = 17 -------------- 63p + 45q = 18 (result of multiplying by 9) 40p - 45q = 85 (result of multiplying by 5) --------------add 103p - 0 = 103 103p = 103 p = 1 now sub 1 in for p in either of the above equations 7p + 5q = 2 7(1) + 5q = 2 7 + 5q = 2 5q = 2 - 7 5q = - 5 q = -1 check... 7p + 5q = 2 7(1) + 5(-1) = 2 7 - 5 = 2 2 = 2 (correct) Now we know that p = 1 and q = -1 -p + 14q -1 + 14(-1) -1 - 14 = - 15

OpenStudy (austinl):

Ah, I see what ya did there jim :D I didn't even realize that. Yeah, just to what jimmy said!

jimthompson5910 (jim_thompson5910):

Subtract the equations 7p + 5q = 2 8p - 9q = 17 ------------ -p + 14q = -15

OpenStudy (texaschic101):

ohhh....thats a lot easier jim

OpenStudy (anonymous):

7p + 5q = 2 -(8p - 9q = 17) gives -p + 14q = - 15

jimthompson5910 (jim_thompson5910):

yes it is, but it only works if everything lines up perfectly (like it does conveniently here) it's probably better to solve for p and q first, then evaluate like you did texaschic101

OpenStudy (texaschic101):

I should have used your way this time....it would have saved time

jimthompson5910 (jim_thompson5910):

true, but you showed how to do it if things didn't match up like they did here so it's good to see the more detailed explanation